#3 A firm issues preferred stock with a dividend of $4.83. If the appropriate discount rate...

#3
A firm issues preferred stock with a dividend of $4.83. If the appropriate discount rate is 9.05% what is the value of the preferred stock?


Submit
Answer format: Currency: Round to: 2 decimal places.


unanswered
not_submitted
#4
A firm will pay a dividend of $1.27 next year. The dividend is expected to grow at a constant rate of 3.12% forever and the required rate of return is 12.30%. What is the value of the stock?


Submit
Answer format: Currency: Round to: 2 decimal places
Answer #1

Q-1

Value of preferred stock =Dividend/discount rate =4.83/9.05% =53.37

Q-2

Value of the stock =Expected dividend/required rate+growth =1.27/(12.3%-3.12%) =13.83